Meeting 2018 03 20

From Hall A Wiki
Jump to: navigation, search

GMn GEM commissioning Meetings: Tuesday 9:00am - 10:00pm

Link to all previous meetings: GMn GEM Commissioning Meeting

Bogdan, Evaristo, Kondo, Leo, Nilanga, Roberto, Siyu

Status of the INFN GEM Chambers

Roberto gave a detailed update on the setup of the cosmic stand for the INFN GEM layers

  • 3 chambers fully instrumented on the cosmic stand (Jo, J1 and J2), all 12 modules are holding HV very well, (nominal 4 kV with Ar-CO2 75/25). One of J2 module that previously had issues (See Evaristo's talk at SBS Coll Meeting Feb. 2018) with HV was also fixed (basically the modules had to be cured over time to burn dusts etc ...)
  • FE cards: Mostly all APV25 cards and back planes on the 3 layers (12 modules) have been tested and are working properly except for a few (3 or 4) cards that for now are been masked. One back plane is also not responding properly, at the moment is also masked
  • DAQ: Main issue is with the DAQ / MPD modules / VME crate. A lot of problems running the system in a stable way even at a very low trigger rate. Roberto and Leo experience full buffer issue when taking data. Seems like some slots in the VME crates are causing the problem. 5 out of 12 of these slots with MPDs had been identified as the source of the problem. There seems to be a pattern with the every two consecutive slots of the crate are OK while the following two are unstable. The team is investigating the problem with the help of Alexandre who also works on the CODA codes to try to debug and fixes some related problems
  • Leo and Roberto are trying to identify 8 stables VME slots to be able to operate 8MPDs to fully read out 2 out 3 GEM layers and start taking cosmic data. They would probably have a clear idea of the situation later today. The priority is to be able to start taking cosmic data in a stable fashion so that we can start characterizing the INFN GEM modules. The DAQ issue should be handle in parallel and would require more experts to take a look at (Alexandre, Danning, Paolo etc ... depending of their availability)
  • Gas bottle: Currently one Ar-CO2 75/25 gas bottle in use, we should always have one spare bottle available during the cosmic data taking (see with Bogdan / Mark how to proceed with this). We also have N2 gas bottle available
  • Trigger: Setup is operational, currently less than 1 Hz trigger rate, can be optimized, we should expect a few hertz not a priority for this week. In addition, one of the scintillator bar of the bottom trigger layer is not operational. We will replace it (discuss wit Bogdan this afternoon about this).
  • Plan is to start data taking today or tomorrow and I will learn how run the system while Leo is still at JLab (until April 3) to take over the monitoring of the cosmic data. Evaristo mention that their is a user manual that can be used by non experts as support when testing the chambers

Status of the UVa GEM Chamber

Siyu gives a brief status of all the items still needed to prepare the UVa GEM layer for the cosmic test

  • All the HDMI cables for the readout electronics are available, same for the patch panels, we need to purchase the male to male adapter for the 12 slots back planes cables
  • The large cable trail is already delivered in the test lab and support plastic plate. Purchase all small items (cable ties etc ...); Siyu is taking care of this
  • Low voltage: Siyu is testing / preparing at UVa the box for the low voltage power supply, checking the fan configuration for cooling the voltage regulators etc ... will report on that at next meeting
  • I'll check with Bogdan for HV cables, possibility for the additional space for the UVa GEM cosmic stand


  • Danning is at JLab today for a couple of days to work with Ben on the implementation of the zero suppression code in the firmware at the SSP level, will get in touch with him